# Must Read: A Systematic Survey of Computational Persuasion

### v1

@@ 2.2.1. Persuasive Strategies & Techniques
Table 1. Example persuasive techniques and sentences designed to discourage smoking from the taxonomy proposed by Zeng et al. (Zeng et al., 2024).
@@
Creating Persuasive Strategy Taxonomies. Prior research has developed various persuasive strategy taxonomies to model and analyze persuasion. Wang et al. (2019) proposed 10 strategies under ”Persuasive Appeal” and ”Persuasive Inquiry,” though some, like ”Donation Information,” are task-specific. Chen and Yang (2021) introduced a more generalized taxonomy with eight strategies, while Zeng et al. (2024) defined a comprehensive set of 40 techniques under 13 umbrella strategies, distinguishing between ”Ethical” and ”Unethical” approaches (See Table 1 for examples). The latter highlights morally ambiguous techniques such as deception, which can enhance persuasive effectiveness, as shown by Durmus et al. (2024), who found that the Claude models are more persuasive when generating deceptive arguments. Dimitrov et al. (2021) identified 22 persuasion techniques in memes, applicable to text and images. Pauli et al. (2022) took a different approach, proposing a more unified computational persuasion taxonomy that frames undesired persuasion as the misuse of rhetorical appeals. Kamali et al. (2024) proposed a novel annotation scheme and dataset of persuasive writing strategies in health misinformation, demonstrating that incorporating these strategies as intermediate signals improves both the accuracy and explainability of misinformation detection models. Other works have also proposed their own taxonomies (Chawla et al., 2021a; Da San Martino et al., 2020; Chen et al., 2021; Piskorski et al., 2023). Despite drawing from established social science research, computational persuasion taxonomies remain highly context-dependent. As a result, the field has yet to establish a unified and generalizable framework for persuasive strategies. Furthermore, it remains an open question whether these strategies influence AI as a Persuadee in the same way they affect humans.
@@
Strategy Classification. Once persuasive strategy taxonomies are established, subsequent research has naturally focused on the automatic detection and classification of these strategies. Accurately classifying persuasive strategies is critical for downstream applications such as detecting automated persuasion and improving the modeling and generation of persuasive language. One of the seminal works in computational persuasion by Wang et al. (2019) introduced a hybrid recurrent convolutional neural network (RCNN) for classifying persuasive strategies in dialogue, reporting 74.8% accuracy and 59.6% macro F1. Similarly, Yang et al. (2019) proposed a semi-supervised neural network to identify persuasive strategies in advocacy requests to help predict the persuasiveness of a message; however, they observed limitations in correctly classifying sentence-level strategies. Chen et al. (2021) later reframed the task as sequence labeling, incorporating intra- and inter-speaker dependencies with a Transformer-based network and an extended Conditional Random Field (CRF). However, despite the adoption of advanced architectures, their models underperformed compared to LSTM-based approaches due to data sparsity and difficulty in modeling label dependencies. Addressing these issues, Chawla et al. (2021a) introduced the CaSiNo dataset, which annotated persuasion strategies in negotiation dialogues and employed a multi-task BERT-based framework to improve classification performance and reported F1 and Joint-A, which is the percentage of utterance strategies predicted correctly. These works show that persuasive strategy prediction is not trivial, and it requires better modeling of context, speaker, and long-distance dependencies. Beyond dialogue, persuasive strategy classification has also been applied to written discourse.
@@
Da San Martino et al. (2020) organized a shared task on propaganda detection, focusing on classifying specific persuasion techniques. Their findings revealed that Transformer-based models dominated the competition, yet simpler strategies with shorter text spans, such as ”Loaded Language,” were classified more effectively, highlighting ongoing challenges in classifying longer, more complex persuasive instances. To develop more generalizable models and make use of document-level persuasion labels, Chen and Yang (2021) proposed a hierarchical weakly-supervised latent variable model that predicts persuasive strategies at the sentence level by leveraging both document- and sentence-level information. Their model outperformed existing semi-supervised baselines, demonstrating the potential of hierarchical learning in persuasion classification.
@@ 2.2.2. Modeling Persuasion
Researchers have explored a range of linguistic, structural, and interactional features to understand what makes an argument persuasive. A widely used resource in this domain is the ChangeMyView (CMV) subreddit, where users present a belief along with supporting reasoning, and others attempt to change their opinion. If a commenter succeeds, the original poster (OP) awards a delta, making CMV a naturally labeled and valuable dataset for studying persuasion in online dialogue.
@@
Bayesian approaches to modeling persuasion have also attracted research interest. Dughmi and Xu (2016) studied this within the sender-receiver framework introduced by Kamenica and Gentzkow (2011b), focusing on computing the sender’s optimal signaling strategy given a prior distribution over payoffs. Wojtowicz (2024) established that informational persuasion is NP-hard. More recently, Li et al. (2025) extended Bayesian persuasion to natural language settings by integrating large language models with game-theoretic techniques. Other tangent work has looked at modeling other attributes, such as deception (Addawood et al., 2019) and intentions (Sakurai and Miyao, 2024; Dutt et al., 2020) in persuasive dialogue and text.
@@ 2.3. Computational Persuasion Taxonomy
[…]
verbatim compartment=(, ),
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, text width=2.7cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2 [Computational Persuasion, fill=gray!45, parent [Evaluating Persuasion, for tree=eval [Detecting Persuasion, eval [(Hidey and McKeown, 2018; Pöyhönen et al., 2022; Shmueli-Scheuer et al., 2019; Dimitrov et al., 2021; Tsinganos et al., 2022; Hasanain et al., 2024), eval_work] ] [Argument Persuasiveness, eval [(Toledo et al., 2019; Habernal and Gurevych, 2016; Simpson and Gurevych, 2018; Pauli et al., 2025; Rescala et al., 2024), eval_work] ] [LLM Persuasiveness, eval [Human Evaluation, eval [(Durmus et al., 2024; OpenAI, 2024, 2025; Phuong et al., 2024; Huang and Wang, 2023; Goldstein et al., 2024; Bai et al., 2025; Salvi et al., 2024), eval_work] ] [Automatic Evaluation, eval [(Singh et al., 2024; Breum et al., 2023; Bozdag et al., 2025; OpenAI, 2024, 2025; Zhu et al., 2025; Idziejczak et al., 2025; Pauli et al., 2025), eval_work] ] ] ] [Generating Persuasion, for tree=gen [Methods, gen [Prompting, gen [(Pauli et al., 2025; Durmus et al., 2024; Zeng et al., 2024; Xu et al., 2024a), gen_work] ] [Incorporating External Information, gen [(Salvi et al., 2024; Zhang and Zhou, 2025; Tiwari et al., 2022b, 2023; Cima et al., 2024; Furumai et al., 2024; Karande et al., 2024), gen_work] ] [Finetuning, gen [(Lewis et al., 2017; Liu et al., 2021; Chen et al., 2022; Jin et al., 2024; Singh et al., 2024), gen_work] ] [Reinforcement Learning, gen [(Tiwari et al., 2022a; Mishra et al., 2022; Shi et al., 2021; Hong et al., 2025), gen_work] ] ] [Applications of Persuasion, gen [Negotiation, gen [(Bianchi et al., 2024; Lewis et al., 2017; Keizer et al., 2017; He et al., 2018; Joshi et al., 2021; Chawla et al., 2021b), gen_work] ] [Debate, gen [(Michael et al., 2023; Du et al., 2024), gen_work] ] [Jailbreaking, gen [(Singh et al., 2023; Zeng et al., 2024; Li et al., 2024; Xu et al., 2024a; Bozdag et al., 2025), gen_work] ] ] ] [Safeguarding Persuasion, for tree=safe [Mitigating Unsafe Persuasion, safe [(Burtell and Woodside, 2023; El-Sayed et al., 2024), safe_work] ] [Selective Acceptance of Persuasion, safe [(Dutt et al., 2021; Sharma et al., 2024; Stengel-Eskin et al., 2025), safe_work] ] ] ]
Figure 2. Taxonomy of computational persuasion.
@@ 3. Evaluating Persuasion
Figure 3. This survey categorizes the evaluation of persuasiveness into three main types: (1) evaluation of argument persuasiveness, (2) human evaluation of LLM-generated content, and (3) automatic evaluation of LLM persuasiveness. For argument persuasiveness, models are typically trained on human-annotated or naturally labeled data to assess the persuasive strength of given arguments. For evaluating LLM persuasiveness, two branches of research emerge: one uses human judges to rate AI-generated content or interactions, while the other relies on LLM-based or non-LLM automatic metrics to perform the evaluation.
Persuasion–in both human discourse and interactions with AI models–poses two intertwined challenges: detecting subtle persuasive cues and evaluating the persuasiveness of the content. Recent advances in natural language processing have helped identify nuanced linguistic markers and argument structures that hint at persuasive intent, even when spread across multi-turn conversations or embedded in multimodal content. Still, both detecting and evaluating the persuasiveness of text segments (i.e., arguments, single-turn utterances, or multi-turn conversations) remains challenging due to the subjectivity of the task and the difficulty of standardizing persuasiveness across multiple domains and contexts. In this section, we review the methodologies on how prior works address these challenges in the detection (§ 3.1) and evaluation of persuasion in arguments (§ 3.2) and LLMs (§ 3.3). As shown in Figure 3, we will discuss persuasion evaluation in three categories: (1) evaluation of argument persuasiveness, (2) human evaluation of LLM persuasiveness, and (3) automatic evaluation of LLM persuasiveness. The main datasets or benchmarks covered in this section are listed in Table 2.
@@ 3.1. Detecting Persuasion
Detecting persuasive cues is critical for identifying when influence is exerted, whether by humans or by language models. Recent advances in natural language processing have significantly improved our ability to discern subtle linguistic patterns and argument structures indicative of persuasive intent. In this section, we review emerging techniques for persuasion detection and discuss their potential to promote more transparent and accountable interactions.
Several studies (Hidey and McKeown, 2018; Pöyhönen et al., 2022; Shmueli-Scheuer et al., 2019; Dimitrov et al., 2021) leveraged machine learning approaches, particularly transformer-based models, to identify persuasive intent in diverse scenarios with textual and conversational data. For instance, Pöyhönen et al. (2022) showed success in training a BERT-based classifier using multilingual semi-annotated dialogues from role-playing games to accurately detect persuasive cues across various languages. Additionally, incorporating personality traits of authors and readers into persuasion detection models has demonstrated significant performance improvements. Shmueli-Scheuer et al. (2019) revealed that personality-aware approaches better capture the nuanced dynamics of persuasion, indicating that personalized modeling can substantially enhance predictive accuracy. Recognizing that persuasion frequently occurs within sequential and contextual frameworks, researchers have also explored neural models capable of handling multi-turn conversational data. Hidey and McKeown (2018) highlighted the importance of analyzing semantic frames and discourse relations to understand how argument ordering influences persuasive impact. This sequential perspective underscores the complexity of persuasive interactions, suggesting that modeling argumentative structure and conversational context is crucial.
@@
ChangeMyView (OpenAI, 2024) LLM Human Evaluation
Persuasion Parallel Generation Evaluation (OpenAI, 2024) LLM Human Evaluation
MakeMePay (OpenAI, 2024) LLM Number of Payments
MakeMeSay (OpenAI, 2024) LLM Game Winrate
@@
Table 2. Datasets or frameworks for persuasiveness evaluation. For datasets or benchmarks without a formal name, we use the name of the paper as the listed name.
3.2. Argument Persuasiveness
Argument persuasiveness has been a longstanding and prominent area of research in natural language processing. The goal is to explore how to automatically assess the persuasive strength of a given argument or a collection of arguments. Evaluation can be conducted in two main ways: absolute evaluation, which assigns a persuasiveness score to individual arguments, and comparative or relative evaluation, which involves ranking or selecting the more persuasive argument from a pair or a group.
A traditional approach to evaluating persuasiveness involves training a specialized model. This typically follows three main steps. First, data is collected through human annotations, which may take the form of either explicit persuasiveness scores assigned to individual texts or preference judgments, where annotators select the more persuasive argument from a pair. Second, the modeling objective is defined. Studies have framed the task as both a pairwise ranking problem, where the model is trained to identify the more persuasive argument from a pair of candidates and as a regression task, aiming to predict a numerical score that reflects the persuasiveness of a given text (Habernal and Gurevych, 2016; Simpson and Gurevych, 2018; Toledo et al., 2019; Pauli et al., 2025). When it is formulated as a pairwise ranking problem, the conventional metric is the accuracy of getting the pairwise ranking correct. On the other hand, when it is a regression task, usually Pearson’s $r$ and Spearman’s $\rho$ are adopted to examine the correlation between the predictions and the ground-truth scores. Third, the model is trained using these objectives. Early work experimented with bidirectional LSTM architectures (Habernal and Gurevych, 2016; Simpson and Gurevych, 2018), while more recent studies have adopted transformer-based models (Toledo et al., 2019; Pauli et al., 2025), which have since become the standard in natural language processing. In the scenario of argument persuasiveness evaluation, transformer-based encoder-only models, like Bert (Devlin et al., 2019), encode text pieces into vectors, which can be further used to perform pairwise ranking tasks or scoring tasks. Whereas decoder-only models, like GPT (Brown et al., 2020; OpenAI, 2022; OpenAI et al., 2024; OpenAI, 2025) are capable of generating utterances as persuaders or persuadees or reason as judges, which is useful in the scenarios described in § 3.3. Saenger et al. (2024) proposed AutoPersuade, a framework for evaluating persuasion through human judgments and a supervised topic model that links argument features to persuasiveness scores. Apart from plain evaluation, AutoPersuade explains what makes arguments more persuasive by estimating the causal effect of interpretable latent topics derived from argument embeddings.

### v2

@@ 2.2. Computational Modeling of Persuasion
Table 1. Sample persuasive techniques designed to discourage smoking from the taxonomy proposed by Zeng et al. (Zeng et al., 2024).
2.2.1. Persuasive Strategies & Techniques
@@
Creating Persuasive Strategy Taxonomies. Prior research has developed various persuasive strategy taxonomies to model and analyze persuasion. Wang et al. (2019) proposed 10 strategies under ”Persuasive Appeal” and ”Persuasive Inquiry,” though some, like ”Donation Information,” are task-specific. Chen and Yang (2021) introduced a more generalized taxonomy with eight strategies, while Zeng et al. (2024) defined a comprehensive set of 40 techniques under 13 umbrella strategies, distinguishing between ”Ethical” and ”Unethical” approaches (See Table 1 for examples). The latter highlights morally ambiguous techniques such as deception, which can enhance persuasive effectiveness Durmus et al. (2024). Dimitrov et al. (2021) identified 22 persuasion techniques in memes, applicable to text and images. Pauli et al. (2022) took a different approach, proposing a more unified computational persuasion taxonomy that frames undesired persuasion as the misuse of rhetorical appeals. Kamali et al. (2024) proposed a novel annotation scheme and dataset of persuasive writing strategies in health misinformation, demonstrating that incorporating these strategies as intermediate signals improves both the accuracy and explainability of misinformation detection models. Other works have also proposed their own taxonomies (Chawla et al., 2021b; Da San Martino et al., 2020; Chen et al., 2021; Piskorski et al., 2023). Despite drawing from established social science research, computational persuasion taxonomies remain highly context-dependent. As a result, the field has yet to establish a unified and generalizable framework for persuasive strategies. Furthermore, it remains an open question whether these strategies influence AI as a Persuadee in the same way they affect humans.
@@
Strategy Classification. Once persuasive strategy taxonomies are established, subsequent research has naturally focused on the automatic detection and classification of these strategies. Accurately classifying persuasive strategies is critical for downstream applications such as detecting persuasion and improving the modeling and generation of persuasive language. One of the seminal works in computational persuasion by Wang et al. (2019) introduced a hybrid recurrent convolutional neural network (RCNN) for classifying persuasive strategies in dialogue. Similarly, Yang et al. (2019) proposed a semi-supervised neural network to identify persuasive strategies in advocacy requests to help predict the persuasiveness of a message; however, they observed limitations in correctly classifying sentence-level strategies. Chen et al. (2021) later reframed the task as sequence labeling, incorporating intra- and inter-speaker dependencies with a Transformer-based network and an extended Conditional Random Field (CRF). However, despite the adoption of advanced architectures, their models underperformed compared to LSTM-based approaches due to data sparsity and difficulty in modeling label dependencies. Addressing these issues, Chawla et al. (2021b) introduced the CaSiNo dataset, which annotated persuasion strategies in negotiation dialogues and employed a multi-task BERT-based framework to improve classification performance. These works show that persuasive strategy prediction is not trivial, and it requires better modeling of context, speaker, and long-distance dependencies. Beyond dialogue, persuasive strategy classification has also been applied to written discourse.
@@
Da San Martino et al. (2020) organized a shared task on propaganda detection, focusing on classifying specific persuasion techniques. Their findings revealed that Transformer-based models dominated the competition, yet simpler strategies with shorter text spans were classified more effectively, highlighting ongoing challenges in classifying longer, more complex persuasive instances. To develop more generalizable models and make use of document-level persuasion labels, Chen and Yang (2021) proposed a hierarchical weakly-supervised latent variable model that predicts persuasive strategies at the sentence level by leveraging both document- and sentence-level information. Their model outperformed existing semi-supervised baselines, demonstrating the potential of hierarchical learning in persuasion classification.
@@ 2.2.2. Modeling Persuasion
Researchers have explored a range of linguistic, structural, and interactional features to understand what makes an argument persuasive. Broadly, persuasion has been modeled through two complementary perspectives: linguistic operationalizations, focusing on textual and stylistic properties that enhance persuasive appeal, and argumentative operationalizations, emphasizing the reasoning structure, discourse relations, and interactional dynamics of persuasion (Bassi et al., 2024). A widely used resource in this domain is the ChangeMyView (CMV) subreddit, where users present a belief along with supporting reasoning, and others attempt to change their opinion. If a commenter succeeds, the original poster (OP) awards a delta, making CMV a naturally labeled and valuable dataset for studying persuasion in online dialogue.
@@
Bayesian approaches to modeling persuasion have also attracted research interest. Dughmi and Xu (2016) studied this within the sender-receiver framework introduced by Kamenica and Gentzkow (2011a), focusing on computing the sender’s optimal signaling strategy given a prior distribution over payoffs. Wojtowicz (2024) established that informational persuasion—the process of selecting and framing subsets of available information to increase a receiver’s belief in a target claim—is NP-hard. More recently, Li et al. (2025) extended Bayesian persuasion to natural language settings by integrating large language models with game-theoretic techniques. Other tangent work has looked at modeling other attributes, such as deception (Addawood et al., 2019) and intentions (Sakurai and Miyao, 2024; Dutt et al., 2020) in persuasive dialogue and text.
@@ 3. Evaluating Persuasion
Persuasion–in both human discourse and interactions with AI models–poses two intertwined challenges: detecting subtle persuasive cues and evaluating the persuasiveness of the content. While recent NLP techniques have begun to explore linguistic markers and argument structures associated with persuasion, reliably identifying such cues across diverse contexts remains difficult. Both detecting and evaluating the persuasiveness of text segments (i.e., arguments, single-turn utterances, or multi-turn conversations) continue to be open challenges due to the subjectivity of the task and the difficulty of standardizing persuasiveness across multiple domains and contexts. In this section, we review the methodologies on how prior works address these challenges in the detection (§ 3.1) and evaluation of persuasion in arguments (§ 3.2) and LLMs (§ 3.3). As shown in Figure 3, we will discuss persuasion evaluation in three categories: (1) evaluation of argument persuasiveness, (2) human evaluation of LLM persuasiveness, and (3) automatic evaluation of LLM persuasiveness. The main datasets or benchmarks covered in this section are listed in Table 2.
@@ 3.1. Detecting Persuasion
Detecting persuasive cues is critical for identifying when influence is exerted, whether by humans or by language models. In this section, we review emerging techniques for persuasion detection and discuss their potential to promote more transparent and accountable interactions. Several studies (Hidey and McKeown, 2018; pöyhönen2022multilingualpersuasiondetectionvideo; Shmueli-Scheuer et al., 2019; Dimitrov et al., 2021) leveraged machine learning approaches, particularly transformer-based models, to identify persuasive intent in diverse scenarios with textual and conversational data. For instance, pöyhönen2022multilingualpersuasiondetectionvideo showed success in training a BERT-based classifier using multilingual semi-annotated dialogues from role-playing games to accurately detect persuasive cues across various languages. Additionally, incorporating personality traits of authors and readers into persuasion detection models has demonstrated significant performance improvements. Shmueli-Scheuer et al. (2019) revealed that personality-aware approaches better capture the nuanced dynamics of persuasion, indicating that personalized modeling can substantially enhance predictive accuracy. Recognizing that persuasion frequently occurs within sequential and contextual frameworks, researchers have also explored neural models capable of handling multi-turn conversational data. Hidey and McKeown (2018) highlighted the importance of analyzing semantic frames and discourse relations to understand how argument ordering influences persuasive impact. This sequential perspective underscores the complexity of persuasive interactions, suggesting that modeling argumentative structure and conversational context is crucial.
@@
Figure 3. This survey categorizes the evaluation of persuasiveness into three main types: (1) evaluation of argument persuasiveness, (2) human evaluation of LLM-generated content, and (3) automatic evaluation of LLM persuasiveness. For argument persuasiveness, models are typically trained on human-annotated or naturally labeled data to assess the persuasive strength of given arguments. For evaluating LLM persuasiveness, two branches of research emerge: one uses human judges to rate AI-generated content or interactions, while the other relies on LLM-based or non-LLM automatic metrics to perform the evaluation.2
3.2. Argument Persuasiveness
Argument persuasiveness has been a longstanding and prominent area of research in NLP. The goal is to learn how to automatically assess the persuasive strength of an argument. Evaluation can be conducted in two main ways: absolute evaluation, which assigns a persuasiveness score to individual arguments, and comparative or relative evaluation, which involves ranking or selecting the more persuasive argument from a pair or a group. Early studies of argument persuasiveness emerged from the broader field of argument mining, which focused on identifying and structuring argumentative components such as claims and premises before progressing toward evaluating their persuasive strength (Lawrence and Reed, 2019; Nguyen and Litman, 2018; Dusmanu et al., 2017; Stab and Gurevych, 2017). This line of research aligns with the broader vision of artificial argumentation, which aims to model how humans construct, exchange, and evaluate arguments, a perspective extensively discussed by Atkinson et al. (2017).
A traditional approach to evaluating persuasiveness involves training a specialized model. This typically follows three main steps. First, data is collected through human annotations, which may take the form of either explicit persuasiveness scores assigned to individual texts or preference judgments, where annotators select the more persuasive argument from a pair. Second, the modeling objective is defined. Studies have framed the task as both a pairwise ranking problem, where the model is trained to identify the more persuasive argument from a pair of candidates and as a regression task, aiming to predict a numerical score that reflects the persuasiveness of a given text (Habernal and Gurevych, 2016; Simpson and Gurevych, 2018; Toledo et al., 2019; Pauli et al., 2025). When it is formulated as a pairwise ranking problem, the conventional metric is the accuracy of getting the pairwise ranking correct. On the other hand, when it is a regression task, usually Pearson’s $r$ and Spearman’s $\rho$ are adopted to examine the correlation between the predictions and the ground-truth scores. Third, the model is trained using these objectives. Early work experimented with bidirectional LSTM architectures (Habernal and Gurevych, 2016; Simpson and Gurevych, 2018), while more recent studies have adopted transformer-based models (Toledo et al., 2019; Pauli et al., 2025), which have since become the standard in natural language processing. In the scenario of argument persuasiveness evaluation, transformer-based encoder-only models, like Bert (Devlin et al., 2019), encode text pieces into vectors, which can be further used to perform pairwise ranking tasks or scoring tasks. Whereas decoder-only models, like GPT (Brown et al., 2020; OpenAI, 2022, 2024a, 2025) are capable of generating utterances as persuaders or persuadees or reason as judges, which is useful in the scenarios described in § 3.3. Saenger et al. (2024) proposed AutoPersuade, a framework for evaluating persuasion through human judgments and a supervised topic model that links argument features to persuasiveness scores. Apart from plain evaluation, AutoPersuade explains what makes arguments more persuasive by estimating the causal effect of interpretable latent topics derived from argument embeddings.
@@ 3.2. Argument Persuasiveness
IBMRank-30k (Gretz et al., 2020) Argument Ranking Scoring (Pointwise Quality)
@@
ChangeMyView (OpenAI, 2024b) LLM Human Evaluation
Persuasion Parallel Generation Evaluation (OpenAI, 2024b) LLM Human Evaluation
MakeMePay (OpenAI, 2024b) LLM Number of Payments
MakeMeSay (OpenAI, 2024b) LLM Game Winrate
@@
Table 2. Datasets or frameworks for persuasiveness evaluation. For those without a formal name, the name of the paper is listed.